# Ill-posedness issue on the Oldroyd-B model in the critical Besov spaces

### v1

@@
Abstract: It is proved in [1, J. Funct. Anal., 2020] that the Cauchy problem for some Oldroyd-B model is well-posed in $\dot{B}^{d/p-1}_{p,1}(\mathbb{R}^{d})\times\dot{B}^{d/p}_{p,1}(\mathbb{R}^{d})$ with $1\leq p<2d$. In this paper, we prove that the Cauchy problem for the same Oldroyd-B model is ill-posed in $\dot{B}^{d/p-1}_{p,r}(\mathbb{R}^{d})\times\dot{B}^{d/p}_{p,r}(\mathbb{R}^{d})$ with $1\leq p\leq\infty$ and $1<r\leq\infty$ due to the lack of continuous dependence of the solution.
@@ 1 Introduction
$\begin{cases}\partial_{t}u+u\cdot\nabla u-\Delta u+\nabla\mathrm{p}=\mu\mathrm{div}\ \tau,&\mathbb{R}^{+}\times\mathbb{R}^{d},\\ \partial_{t}\tau+u\cdot\nabla\tau+\tau\omega-\omega\tau=0,&\mathbb{R}^{+}\times\mathbb{R}^{d},\\ \mathrm{div}\ u=0,&\mathbb{R}^{+}\times\mathbb{R}^{d},\\ (u,\tau)(t=0)=\left(u_{0},\tau_{0}\right),&\mathbb{R}^{d}.\end{cases}$ (1.1)
Here the parameter $\mu\geq 0$. The unknown function $u=u(t,x)$ is the velocity field of a particle $x\in\mathbb{R}^{n}$ at a time $t\in\mathbb{R}$, $\tau=(\tau(t,x))_{ij}$ with $1\leq i,j\leq d$ is the conformation tensor in $\mathbb{R}^{d\times d}$, which describes the internal elastic forces that the constitutive molecules exert on each other, and $\mathrm{p}$ is the pressure. The evolutionary equation for the conformation tensor $\tau$ is then driven by the vorticity tensor $\omega$, which stands for the skew-adjoint part of the deformation tensor $\nabla u$:
@@
$\begin{cases}\partial_{t}u+u\cdot\nabla u-\nu\Delta u+\nabla\mathrm{p}=\mu_{1}\operatorname{div}\tau,&\mathbb{R}^{+}\times\mathbb{R}^{d},\\ \partial_{t}\tau+u\cdot\nabla\tau+\alpha\tau+\mathbb{Q}(u,\tau)=\mu_{2}\mathbb{D}(u),&\mathbb{R}^{+}\times\mathbb{R}^{d},\\ \operatorname{div}u=0,&\mathbb{R}^{+}\times\mathbb{R}^{d},\\ (u,\tau)(t=0)=\left(u_{0},\tau_{0}\right),&\mathbb{R}^{d}.\end{cases}$ (1.2)
@@
$\mu_{1}=\nu=1,\quad\mu_{2}=\mu\geq 0,\quad\alpha=\beta=0,$ (1.3)
@@ 1.1 A review of related results
$\left(u_{\ell}(t,x),\tau_{\ell}(t,x)\right)=\left(\ell u(\ell^{2}t,\ell x),\ell^{2}\tau(\ell^{2}t,\ell x)\right),\quad\ell>0.$
This suggests us to choose initial data $(u_{0},\tau_{0})$ in “critical spaces” whose norm is invariant for all $\ell>0$ (up to a constant independent of $\ell$) by the transformation $(u_{0},\tau_{0})(x)\mapsto(\ell u_{0}(\ell x),\ell^{2}\tau_{0}(\ell x))$. It is natural that $\dot{B}_{p,r}^{{d}/{p}-1}\times\dot{B}_{p,r}^{{d}/{p}}$ are critical spaces to the Oldroyd-B model (1.1). Chemin and Masmoudi [4] first established the global well-posedness result of system (1.2) in the critical $L^{p}$ framework. They particularly showed the local and global-in-time existence of solutions for large and small initial data, respectively, under the assumption of a smallness condition on the coupling parameters of system (1.2). Some improvements were made by Chen and Miao in [6]. Zi, Fang and Zhang [29] removed the smallness restriction on the coupling parameter. Fang and Zi [14] also constructed strong solutions to the Oldroyd-B model (1.2) with large vertical initial velocity in the critical $L^{2}$ framework. For more results on the Oldroyd-B model, we refer to see [5, 19, 11, 25, 26, 28] and the references therein.
Anna and Paicu [1] studied the Oldroyd-B model (1.1) under suitable condition on the initial data and they showed:
  • existence of global-in-time classical solutions in dimension two for large initial data,
  • uniqueness in dimension two of strong solutions,
  • existence and uniqueness of global-in-time strong solution in dimension $d\geq 3$, with a Fujita-Kato smallness condition for the initial data.
In particular, they established the following well-posedness result (see [1, Theorem 1.5]).
Theorem 1.1.
Let $p\in[1,2d)$ and $d\geq 3$. Assume that the initial data $u_{0}$ be a free divergence vector field in $\dot{B}_{p,1}^{\frac{d}{p}-1}$, while $\tau_{0}$ belongs to $\dot{B}_{p,1}^{\frac{d}{p}}$. Then there exists a time $T^{*}>0$ for which system (1) admits a unique local solution $(u,\tau)$ satisfying
@@
From the PDE’s point of view, it is crucial to know if an equation which models a physical phenomenon is well-posed in the Hadamard’s sense: existence, uniqueness, and continuous dependence of the solutions with respect to the initial data. In particular, the lack of continuous dependence would cause incorrect solutions or non meaningful solutions. Indeed, this means that the corresponding equation is ill-posed. Many results with regard to the ill-posedness have been obtained for some important nonlinear PDEs including the incompressible Navier-Stokes equations [3, 24, 27], the compressible Navier-Stokes equations [7, 9, 17] and the incompressible non-resistive MHD equations [8] and so on.
In this paper, we are mainly focused on the ill-posedness of the Cauchy problem (1.1) in some critical Besov spaces. Precisely speaking, for the case $1<r\leq\infty$, it is still unknown whether the Cauchy problem (1.1) in $\dot{B}^{\frac{d}{p}-1}_{p,r}(\mathbb{R}^{d})\times\dot{B}^{\frac{d}{p}}_{p,r}(\mathbb{R}^{d})$ is well-posed or ill-posed. In this paper, we shall answer this question.
@@ 1.2 Main Result
Let $d\geq 2$, $1\leq p\leq\infty$ and $1<r\leq\infty$. The Cauchy problem (1.1) is ill-posed in $\dot{B}^{\frac{d}{p}-1}_{p,r}(\mathbb{R}^{d})\times\dot{B}^{\frac{d}{p}}_{p,r}(\mathbb{R}^{d})$. More precisely, there exist a sequence of initial data $(u^{n}_{0},\tau^{n}_{0})$, positive time $\left\{T_{n}\right\}_{n}$ with $T_{n}\rightarrow 0$ as $n\rightarrow\infty$, and a sequence of corresponding smooth solutions $\left(u^{n},\tau^{n}\right)\in\mathcal{C}([0,T_{n}];\dot{B}^{\frac{d}{p}-1}_{p,r})\times\mathcal{C}([0,T_{n}];\dot{B}^{\frac{d}{p}}_{p,r})$ such that
@@
$\|\tau^{n}(T_{n},\cdot)\|_{\dot{B}^{\frac{d}{p}}_{p,r}}\geq c_{0}$
@@
Theorem 1.2 demonstrates that if $d\geq 2$, $1\leq p\leq\infty$ and $1<r\leq\infty$, there exists a sequence of initial data which converges to zero in $\dot{B}^{\frac{d}{p}-1}_{p,r}(\mathbb{R}^{d})\times\dot{B}^{\frac{d}{p}}_{p,r}(\mathbb{R}^{d})$ and yields a sequence of solutions to (1.1) which does not converge to zero in $\dot{B}^{\frac{d}{p}-1}_{p,r}(\mathbb{R}^{d})\times\dot{B}^{\frac{d}{p}}_{p,r}(\mathbb{R}^{d})$. In other words, (1.1) is ill-posed in $\dot{B}^{\frac{d}{p}-1}_{p,r}(\mathbb{R}^{d})\times\dot{B}^{\frac{d}{p}}_{p,r}(\mathbb{R}^{d})$ due to the discontinuity of the solution map at zero.
@@ 1.3 Main Idea
We try to extract the worst nonlinear term $P$ from the $\tau$-equation, which leads to the discontinuous solution map of $\tau$ at time $t=0$ in the metric of $\dot{B}^{\frac{d}{p}}_{p,r}$.
@@ 2 Preliminaries
The inverse Fourier transform of any $g$ is given by
$(\mathcal{F}^{-1}g)(x)=\check{g}(x)=\frac{1}{(2\pi)^{d}}\int_{\mathbb{R}^{d}}g(\xi)e^{\mathrm{i}x\cdot\xi}\mathrm{d}\xi.$
@@
For every $u\in\mathcal{S^{\prime}}(\mathbb{R}^{d})$, the homogeneous dyadic blocks ${\dot{\Delta}}_{j}$ is defined as follows
$\dot{\Delta}_{j}u=\varphi(2^{-j}D)u=\mathcal{F}^{-1}\big{(}\varphi(2^{-j}\cdot)\mathcal{F}u\big{)}=2^{dj}\int_{\mathbb{R}^{d}}\check{\varphi}\big{(}2^{j}(x-y)\big{)}u(y)\mathrm{d}y,\quad\forall j\in\mathbb{Z}.$
@@
$\mathcal{S}^{\prime}_{h}:=\Big{\{}u\in\mathcal{S^{\prime}}(\mathbb{R}^{d}):\;\lim_{j\rightarrow-\infty}\|\chi(2^{-j}D)u\|_{L^{\infty}}=0\Big{\}}.$
@@
Let $1\leq\rho,\rho_{1},\rho_{2}\leq\infty$ with $\frac{1}{\rho}=\frac{1}{\rho}_{1}+\frac{1}{\rho}_{2}$ and $1\leq p<2d$. Then, we have
@@
$\|u\|_{\tilde{L}_{T}^{\infty}\left(\dot{B}_{p,1}^{s}\right)}\leq e^{CV_{p_{1}}(t)}\left(\left\|u_{0}\right\|_{\dot{B}_{p,1}^{s}}+\int_{0}^{t}e^{-CV_{p_{1}}(\tau)}\|g(\tau)\|_{\dot{B}_{p,1}^{s}}\mathrm{~{}d}\tau\right),$
where $V_{p_{1}}(t)=\int_{0}^{t}\|\nabla v\|_{\dot{B}_{p_{1},1}^{p_{1}}\left(\mathbb{R}^{d}\right)}\mathrm{d}s$.
@@
$\left\{\begin{array}[]{lr}\partial_{t}u-\nu\Delta u=f\in\tilde{L}^{\rho_{1}}\left(0,T;\dot{B}_{p,r}^{s-2+2/\rho_{1}}(\mathbb{R}^{d})\right),\\ u(t=0)=u_{0}\in\dot{B}_{p,r}^{s}(\mathbb{R}^{d}),\end{array}\right.$
@@
$\nu^{\frac{1}{\rho}}\|u\|_{\tilde{L}^{\rho}_{t}\left(\dot{B}_{p,r}^{s+\frac{2}{\rho}}\right)}\leq C\left(\left\|u_{0}\right\|_{\dot{B}_{p,r}^{s}}+\nu^{\frac{1}{\rho_{1}}-1}\|f\|_{\tilde{L}^{\rho_{1}}_{t}(\dot{B}_{p,r}^{s-2+\frac{2}{\rho_{1}}})}\right),\quad\rho\geq\rho_{1}.$
@@
Assume that $u_{0}\in\dot{B}_{p,r}^{s}$ and $f\in\tilde{L}^{1}\left(0,T;\dot{B}_{p,r}^{s}\right)$, then there exists a unique solution $u$ satisfying
@@
$\nu^{\frac{1}{\rho}}\|u\|_{\tilde{L}^{\rho}\left(0,T;\dot{B}_{p,r}^{s+\frac{2}{\rho}}\right)}\leq C\left(\left\|u_{0}\right\|_{\dot{B}_{p,r}^{s}}+\|\mathcal{P}f\|_{L^{1}\left(0,T;\dot{B}_{p,r}^{s}\right)}\right).$
@@
Assume that $u$ is a smooth vector field and $\phi(t,x)$ satisfies (1.5). Then, for all $t\in\mathbb{R}^{+}$, the flow $\phi(t,x)$ is a $C^{1}$ diffeomorphism over $\mathbb{R}^{d}$, and we have
@@
We prove Theorem 1.2 by dividing it into several parts: (1) Construction of initial data; (2) Estimation of initial data; (3) Local well-posedness for (1.1) with initial data (3.9)-(3.10); (4) Discontinuity of the solution map.

### v2

@@
Abstract: It was proved in [1, J. Funct. Anal., 2020] that the Cauchy problem for some Oldroyd-B model is well-posed in $\dot{B}^{d/p-1}_{p,1}(\mathbb{R}^{d})\times\dot{B}^{d/p}_{p,1}(\mathbb{R}^{d})$ with $1\leq p<2d$. In this paper, we prove that the Cauchy problem for the same Oldroyd-B model is ill-posed in $\dot{B}^{d/p-1}_{p,r}(\mathbb{R}^{d})\times\dot{B}^{d/p}_{p,r}(\mathbb{R}^{d})$ with $1\leq p\leq\infty$ and $1<r\leq\infty$ due to the lack of continuous dependence of the solution.
@@ 1 Introduction
$\begin{cases}\partial_{t}u+u\cdot\nabla u-\Delta u+\nabla\mathrm{p}=\mathrm{div}\ \tau,&(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{d},\\ \partial_{t}\tau+u\cdot\nabla\tau+\tau\omega-\omega\tau=0,&(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{d},\\ \mathrm{div}\ u=0,&(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{d},\\ (u,\tau)(t=0)=\left(u_{0},\tau_{0}\right),&x\in\mathbb{R}^{d},\end{cases}$ (1.1)
The unknown function $u=u(t,x)$ is the velocity field of a particle $x\in\mathbb{R}^{d}$ at a time $t\in\mathbb{R}^{+}$, $\tau=(\tau(t,x))_{ij}$ with $1\leq i,j\leq d$ is the conformation tensor in $\mathbb{R}^{d\times d}$, which describes the internal elastic forces that the constitutive molecules exert on each other, and $\mathrm{p}(t,x)$ is the pressure. The evolutionary equation for the conformation tensor $\tau$ is then driven by the vorticity tensor $\omega$, which stands for the skew-adjoint part of the deformation tensor $\nabla u$:
@@
$\begin{cases}\partial_{t}u+u\cdot\nabla u-\nu\Delta u+\nabla\mathrm{p}=\mu_{1}\operatorname{div}\tau,&(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{d},\\ \partial_{t}\tau+u\cdot\nabla\tau+\alpha\tau+\mathbb{Q}(u,\tau)=\mu_{2}\mathbb{D}(u),&(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{d},\\ \operatorname{div}u=0,&(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{d},\\ (u,\tau)(t=0)=\left(u_{0},\tau_{0}\right),&x\in\mathbb{R}^{d}.\end{cases}$ (1.2)
@@
$\mu_{1}=\nu=1,\quad\mu_{2}=0,\quad\alpha=\beta=0,$ (1.3)
@@ 1.1 A review of related results
$\left(u_{\ell}(t,x),\tau_{\ell}(t,x)\right)=\left(\ell u(\ell^{2}t,\ell x),\tau(\ell^{2}t,\ell x)\right),\quad\ell>0.$
This suggests us to choose initial data $(u_{0},\tau_{0})$ in “critical spaces” whose norm is invariant for all $\ell>0$ (up to a constant independent of $\ell$) by the transformation $(u_{0},\tau_{0})(x)\mapsto(\ell u_{0}(\ell x),\tau_{0}(\ell x))$. It is natural that $\dot{B}_{p,r}^{{d}/{p}-1}\times\dot{B}_{p,r}^{{d}/{p}}$ are critical spaces to the Oldroyd-B model (1.1). Chemin and Masmoudi [4] first established the global well-posedness result of system (1.2) in the critical $L^{p}$ framework. They particularly showed the local and global-in-time existence of solutions for large and small initial data, respectively, under the assumption of a smallness condition on the coupling parameters of system (1.2). Some improvements were made by Chen and Miao in [6]. Zi, Fang and Zhang [29] removed the smallness restriction on the coupling parameter. Fang and Zi [14] also constructed strong solutions to the Oldroyd-B model (1.2) with large vertical initial velocity in the critical $L^{2}$ framework. For more results on the Oldroyd-B model, we refer to see [5, 19, 11, 25, 26, 28] and the references therein.
Anna and Paicu [1] studied the Oldroyd-B model (1.1) under suitable condition on the initial data and they showed the existence of global-in-time classical solutions in dimension two for large initial data, uniqueness in dimension two of strong solutions, and existence and uniqueness of global-in-time strong solution in dimension $d\geq 3$ with a Fujita-Kato smallness condition for the initial data. In particular, they established the following well-posedness result
Theorem 1.1 ([1]).
Let $p\in[1,2d)$ and $d\geq 3$. Assume that the initial data $u_{0}$ be a free divergence vector field in $\dot{B}_{p,1}^{\frac{d}{p}-1}$, while $\tau_{0}$ belongs to $\dot{B}_{p,1}^{\frac{d}{p}}$. Then there exists a time $T^{*}>0$ for which system (1.1) admits a unique local solution $(u,\tau)$ satisfying
@@
From the PDE’s point of view, it is crucial to know if an equation which models a physical phenomenon is well-posed in the Hadamard’s sense: existence, uniqueness, and continuous dependence of the solutions with respect to the initial data. In particular, the lack of continuous dependence would cause incorrect solutions or non meaningful solutions. Indeed, this means that the corresponding equation is ill-posed. Many results with regard to the ill-posedness for some important nonlinear PDEs have been obtained. We would like to mention the following incompressible non-resistive magnetohydrodynamics (MHD) equations
$\begin{cases}\partial_{t}u+u\cdot\nabla u-\Delta u+\nabla\mathrm{p}=b\cdot\nabla b,&(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{d},\\ \partial_{t}b+u\cdot\nabla b=b\cdot\nabla u,&(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{d},\\ \mathord{{\rm div}}\,u=\mathord{{\rm div}}\,b=0,&(t,x)\in\mathbb{R}^{+}\times\mathbb{R}^{d},\\ (u,b)(t=0)=(u_{0},b_{0}),&x\in\mathbb{R}^{d}.\end{cases}$ (1.4)
Here the initial data $(u_{0},b_{0})$ is divergence-free. $u=u(t,x)$ represents the velocity field, $b=b(t,x)$ denotes the magnetic field and $\mathrm{p}=\mathrm{p}(t,x)$ is the scalar pressure, respectively. Chen, Nie and Ye [8] proved that the ill-posedness of the d-dimension $(d\geq 2)$ MHD equations (1.4) in $\dot{B}^{d/p-1}_{p,r}(\mathbb{R}^{d})\times\dot{B}^{d/p}_{p,r}(\mathbb{R}^{d})$ with $1\leq p\leq\infty$ and $r>1$ by constructing complex-valued initial data for showing the norm inflation of the magnetic field which comes from the analysis of nonlinear term $b\cdot\nabla u$. We refer to see more ill-posedness results on the incompressible Navier-Stokes equations [3, 24, 27], the compressible Navier-Stokes equations [7, 9, 17] and so on. In this paper, we are mainly focused on the ill-posedness of the Cauchy problem (1.1) in some critical Besov spaces. Precisely speaking, for the case $1<r\leq\infty$, it is still unknown whether the Cauchy problem (1.1) in $\dot{B}^{d/p-1}_{p,r}(\mathbb{R}^{d})\times\dot{B}^{d/p}_{p,r}(\mathbb{R}^{d})$ is well-posed or ill-posed. In this paper, we shall answer this question.
@@ 1.2 Main Result
Let $d\geq 2$, $1\leq p\leq\infty$ and $1<r\leq\infty$. The Cauchy problem (1.1) is ill-posed in $\dot{B}^{d/p-1}_{p,r}(\mathbb{R}^{d})\times\dot{B}^{d/p}_{p,r}(\mathbb{R}^{d})$. More precisely, there exist a sequence of initial data $(u^{n}_{0},\tau^{n}_{0})\in\mathcal{S}(\mathbb{R}^{d})$, positive time $\left\{T_{n}\right\}_{n}$ with $T_{n}\rightarrow 0$ as $n\rightarrow\infty$, and a sequence of corresponding smooth solutions $\left(u^{n},\tau^{n}\right)$ such that
@@
$\left\|\tau^{n}(T_{n},\cdot)\right\|_{\dot{B}^{\frac{d}{p}}_{p,r}}\geq c_{0}$
@@
Theorem 1.2 demonstrates that if $d\geq 2$, $1\leq p\leq\infty$ and $1<r\leq\infty$, there exists a sequence of initial data which converges to zero in $\dot{B}^{d/p-1}_{p,r}(\mathbb{R}^{d})\times\dot{B}^{d/p}_{p,r}(\mathbb{R}^{d})$ and yields a sequence of solutions to (1.1) which does not converge to zero in $\dot{B}^{d/p-1}_{p,r}(\mathbb{R}^{d})\times\dot{B}^{d/p}_{p,r}(\mathbb{R}^{d})$. In other words, (1.1) is ill-posed in $\dot{B}^{d/p-1}_{p,r}(\mathbb{R}^{d})\times\dot{B}^{d/p}_{p,r}(\mathbb{R}^{d})$ due to the discontinuity of the solution map at zero.
Remark 1.2.
We would like to mention that, Theorem 1.2 holds for the case when $p=r=2$ and demonstrates that (1.1) is ill-posed in the critical Sobolev spaces $\dot{H}^{d/2-1}(\mathbb{R}^{d})\times\dot{H}^{d/2}(\mathbb{R}^{d})$.
@@ 1.3 Main Idea
We try to extract the worst nonlinear term $P$ from the $\tau$-equation, which leads to the discontinuous solution map of $\tau$ at time $t=0$ in the metric of $\dot{B}^{d/p}_{p,r}$.
@@
Remark 1.3.
We would like to point out the main difference between our contribution and Chen-Nie-Ye’s work [8]. Compared with the initial velocity field $u_{0}$ and initial magnetic field $b_{0}$ which are complex-valued in [8], both initial data $u_{0}$ and $\tau_{0}$ in this work are real-valued and seem to be more physically significant. Also, our construction of $(u_{0},\tau_{0})$ makes the the analysis and computations more clean and may be enlightening to the ill-posedness for some important nonlinear PDEs. In summary, we proved that the Cauchy problem for the Oldroyd-B model (1.1) with real-valued initial data is ill-posed in $\dot{B}^{d/p-1}_{p,r}(\mathbb{R}^{d})\times\dot{B}^{d/p}_{p,r}(\mathbb{R}^{d})$ with $1\leq p\leq\infty$ and $1<r\leq\infty$ while it is not clear that the Cauchy problem for the MHD equations (1.4) with real-valued initial data is ill-posed in the same spaces.
@@ 2 Preliminaries
The inverse Fourier transform allows us to recover $f$ from $\widehat{f}$
$f(x)=\mathcal{F}^{-1}\widehat{f}(x)=(2\pi)^{-d}\int_{\mathbb{R}^{d}}e^{\mathrm{i}x\cdot\xi}\widehat{f}(\xi)\mathrm{d}\xi.$
@@
For every $u\in\mathcal{S^{\prime}}(\mathbb{R}^{d})$, the homogeneous dyadic block ${\dot{\Delta}}_{j}$ is defined as follows
$\dot{\Delta}_{j}u=\varphi(2^{-j}D)u=\mathcal{F}^{-1}\big{(}\varphi(2^{-j}\cdot)\mathcal{F}u\big{)}=2^{dj}\int_{\mathbb{R}^{d}}\check{\varphi}\big{(}2^{j}(\cdot-y)\big{)}u(y)\mathrm{d}y,\quad\forall j\in\mathbb{Z}.$
@@
$\mathcal{S}^{\prime}_{h}:=\Big{\{}u\in\mathcal{S^{\prime}}(\mathbb{R}^{d}):\;\lim_{j\rightarrow-\infty}\left\|\vartheta(2^{-j}D)u\right\|_{L^{\infty}}=0\Big{\}}.$
@@
Let $1\leq\rho,\rho_{1},\rho_{2}\leq\infty$ with $\frac{1}{\rho}=\frac{1}{\rho}_{1}+\frac{1}{\rho}_{2}$ and $1\leq p<2d$. Then
@@
$\|u\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{p,1}^{s})}\leq e^{CV_{p_{1}}(t)}\left(\left\|u_{0}\right\|_{\dot{B}_{p,1}^{s}}+\int_{0}^{t}e^{-CV_{p_{1}}(\tau)}\|g(\tau)\|_{\dot{B}_{p,1}^{s}}\mathrm{~{}d}\tau\right),$
where $V_{p_{1}}(t)=\int_{0}^{t}\|\nabla v\|_{\dot{B}_{p_{1},1}^{p_{1}}(\mathbb{R}^{d})}\mathrm{d}s$.
@@
$\begin{cases}\partial_{t}u-\nu\Delta u=f\in\tilde{L}^{\rho_{1}}\left(0,T;\dot{B}_{p,r}^{s-2+2/\rho_{1}}(\mathbb{R}^{d})\right),\\ u(t=0)=u_{0}\in\dot{B}_{p,r}^{s}(\mathbb{R}^{d}),\end{cases}$
@@
$\nu^{\frac{1}{\rho}}\|u\|_{\tilde{L}^{\rho}_{t}(\dot{B}_{p,r}^{s+\frac{2}{\rho}})}\leq C\left(\left\|u_{0}\right\|_{\dot{B}_{p,r}^{s}}+\nu^{\frac{1}{\rho_{1}}-1}\|f\|_{\tilde{L}^{\rho_{1}}_{t}(\dot{B}_{p,r}^{s-2+\frac{2}{\rho_{1}}})}\right),\quad\rho\geq\rho_{1}.$
@@
Assume that $u_{0}\in\dot{B}_{p,r}^{s}$ and $f\in\tilde{L}^{1}\left(0,T;\dot{B}_{p,r}^{s}\right)$, then there exists a unique solution $(u,\mathrm{p})$ satisfying
@@
$\nu^{\frac{1}{\rho}}\|u\|_{\tilde{L}^{\rho}(0,T;\dot{B}_{p,r}^{s+\frac{2}{\rho}})}\leq C\left(\left\|u_{0}\right\|_{\dot{B}_{p,r}^{s}}+\|\mathcal{P}f\|_{L^{1}\left(0,T;\dot{B}_{p,r}^{s}\right)}\right).$
@@
Assume that $u$ is a smooth vector field and $\phi(t,x)$ satisfies (1.6). Then, for all $t\in\mathbb{R}^{+}$, the flow $\phi(t,x)$ is a $C^{1}$-diffeomorphism over $\mathbb{R}^{d}$, and we have
@@
In this section, we prove Theorem 1.2 by dividing it into several parts: (1) Construction of initial data; (2) Estimation of initial data; (3) Local well-posedness for (1.1) with initial data (3.12)-(3.13); (4) Discontinuity of the solution map.